A skier starting from rest accelerates uniformly downhill at 1.8 m/s2 [fwd]. 

How long will it take the skier to reach a point 95 m [fwd] from the starting position?




1
Expert's answer
2022-01-13T09:21:58-0500

The time "t" can be found from the kinematic formula:


"d = \dfrac{at^2}{2}"

where "d = 95m" and "a = 1.8m\/s^2". Thus, obtain:


"t = \sqrt{\dfrac{2d}{a}} = \sqrt{\dfrac{2\cdot 95m}{1.8m/s^2}} \approx 10s"

Answer. 10s.